OpenStudy (anonymous):

How much money invested at 5% compounded continuously for 3 years will yield $820?

OpenStudy (2yoututu11):

$820=Pe^.15 $820=P*1.16 P=820/1.16=$706.90

OpenStudy (anonymous):

$708.34

OpenStudy (anonymous):

is this the correct answer?? Hope so!!